OpenStudy (anonymous):

Please help will fan and medal! Evaluate the expression by using the given values of the variables. y ÷ 4 • 3 – 6 ÷ x + 8 where x = 3 and y = 24 A) 8 Eliminate B) 12 C) 24 D) –8

OpenStudy (anonymous):

Beforehand thanks for all the help

OpenStudy (anonymous):

Insert the values for x and y. \[24/4*3-6/x+8=?\]

OpenStudy (anonymous):

Meant to write 3 instead of x.

OpenStudy (anonymous):

um can u simplify that for me? lol

OpenStudy (landon34):

you can simplify it using http://www.geteasysolution.com/

OpenStudy (anonymous):

Do you know the order of operations?

OpenStudy (anonymous):

i tried but it says it has to have an equal sign

OpenStudy (anonymous):

@Landon34

OpenStudy (landon34):

did you not put in the equal sign?

OpenStudy (anonymous):

i thought that might change it completely. I am so stupid lol

OpenStudy (landon34):

Ill try putting in the equation.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[3*(24/4)-(6/3)+8 = 0\] \[3*6-(6/3)+8 = 0\] \[8-(6/3)+18 = 0\] \[8-2+18 = 0\] \[8+16 = 0\] \[24 = 0\]

OpenStudy (landon34):

never mind.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Any help ppl?

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (landon34):

he put in the simplified equation up there.

OpenStudy (anonymous):

rlly? I will check it out

OpenStudy (landon34):

no, i mean up there.

OpenStudy (anonymous):

omg i forgot evaluate meant solve hold on i gonna solve it real quick.

OpenStudy (anonymous):

The answer is 12. Thanks @Tommynaiter and @Landon34 !

OpenStudy (anonymous):

but who do i give medal to?

OpenStudy (landon34):

ok, youre welcome bruh.

OpenStudy (anonymous):

How did you get 12?

OpenStudy (landon34):

I can give one to tommynaiter.

OpenStudy (anonymous):

i solved the whole equation.

OpenStudy (anonymous):

You did something wrong, because it is 24. Can I see your work?

OpenStudy (anonymous):

i forgot for a second evaluate meant solve lol.

OpenStudy (anonymous):

yes. here it is- 24/4*3-6/3+8 order of operations Parenthesis Exponents Multiplication Division Add Subtract omg i so wrong!

OpenStudy (anonymous):

i didn't follow order of operation!

OpenStudy (anonymous):

lemme solve again.

OpenStudy (anonymous):

Haha, no worries try and see if you get 24

OpenStudy (landon34):

ok.

OpenStudy (anonymous):

ok i will see

OpenStudy (anonymous):

i got 8 when i followed order of operations....

OpenStudy (anonymous):

lemme guess i did it wrong. AGAIN.:(

OpenStudy (anonymous):

I know it 24 but i wanna know how to get it.

OpenStudy (anonymous):

@Tommynaiter?

OpenStudy (anonymous):

weird

OpenStudy (landon34):

hes gone. :/

OpenStudy (anonymous):

i will give u a medal @Landon34 I guess

OpenStudy (landon34):

the answer is 24, bruh.

OpenStudy (anonymous):

I know it 24 but he left so here the medal

OpenStudy (anonymous):

oh wait he back

OpenStudy (anonymous):

I showed my work, some time ago, go up and see it if you'd like :)

OpenStudy (anonymous):

ok i will here your medal

OpenStudy (anonymous):

OMG I DIDN"T SEE THAT EARLIER!

OpenStudy (anonymous):

XD TYSM @Tommynaiter!

OpenStudy (landon34):

i told you to scroll up bruh, you didnt see my post?

OpenStudy (anonymous):

AND YOUR HELP TOO @Landon34!

OpenStudy (anonymous):

no i didn't.

OpenStudy (landon34):

ok, your welcome @ldhmorris

OpenStudy (anonymous):

lol i am such a klutz i am closing question now bye.